Grade Select GradeAlgebra

3x+y=5 and x+y+1=0 are two diameters to the circle which passes through the point (-2,2).Find its equation .Also find the radius of the circle

aradhya singh
9 Years agoGrade Select Grade

1 Answer

Vikas TU
9 Years ago
since 3x+y=5 and x+y+1=0 are two diameters .Intersection point of these two diameters will be center of the circle.
Solving 3x+y=5 and x+y+1=0 the intersection point will be (3,-4)
Center of the circle is (3,-4)
Radius is the distance between (3,-4) and (-2,2)
Applying distance formula ((3 – (-2))^2 + (-4 -2 )^2)^ ½
Radius is (61)^ ½
Equation of circle is (x-3)^2 + (y+4)^2 = 61.